solve for x.
4x(x-8)-5x(x-7)=-4

Answer by stanbon(75887)

solve for x.
4x(x-8)-5x(x-7)=-4
4x^2-8x-5x^2+35x = -4
------
-x^2+17x = -4
----
x^2-17x-4 = 0
---
x = [17+-sqrt(17^2-4*1*-4)]/2
---
x = [17 +- sqrt(305)]/2
---
x = [17 +- 17.46]/2
----
x = 17.23 or x = -0.23
==============================
Cheers,
Stan H.